OpenStudy (anonymous):

!@# how do i FACTOR this into this?? 2u^2-5u+2=0 to get (2u-1) (u-2)= 0 i forgot the steps to do it??

jimthompson5910 (jim_thompson5910):

2u^2-5u+2=0 2u^2 - 4u - u+2=0 Note: -4 and -1 multiply to 2*2 = 4 and add to -5 (2u^2-4u) + (-u + 2) = 0 2u(u-2) -1(u - 2) = 0 (2u-1)(u-2) = 0 I'll let you take over from here.

OpenStudy (anonymous):

how do i know to change the -5u into -4u-1u

jimthompson5910 (jim_thompson5910):

You multiply the first coefficient by the last term to get: 2*2 = 4 Then you look for two numbers that multiply to that last result AND add to the middle coefficient of -5 These two numbers are -4 and -1 So this means that you break up -5u into -4u - 1u or just -4u - u